OpenStudy (kitten_is_back):

One isotope of carbon has 6 protons and 8 neutrons. Which is the correct reference for this isotope? A. carbon-6 B. carbon-8 C. carbon-14 D. carbon-2

OpenStudy (anonymous):

atomic mass = 6 + 8 = 14 it is 14-C so answer is option c=carbon-14

OpenStudy (kitten_is_back):

thanks

OpenStudy (anonymous):

yw